# A Novel Method to Constrain Tidal Quality Factor from A Non-synchronized Exoplanetary System

### v1

@@ Abstract
We propose a novel method to constrain the tidal quality factor, $Q^{\prime}$, from a non-synchronized star-planet system consisting of a slowly rotating low-mass star and a close-in Jovian planet, taking into account the tidal interaction and the magnetic braking. On the basis of dynamical system theory, the track of the co-evolution of angular momentum for such a system exhibits the existence of a forbidden region in the $\Omega_{\mathrm{orb}}$$\Omega_{\mathrm{spin}}$ plane , where $\Omega_{\mathrm{spin}}$ and $\Omega_{\mathrm{orb}}$ denote the angular velocity of the stellar spin and planetary orbit, respectively. The forbidden region is determined primarily by the strength of the tidal interaction. By comparing ($\Omega_{\mathrm{orb}},\Omega_{\mathrm{spin}}$) of a single star-planet system to the forbidden region, we can constrain the tidal quality factor regardless of the evolutionary history of the system. The application of this method to the star-planet system, NGTS-10 – NGTS-10 b, gives $Q^{\prime}\gtrsim 10^{8}$, leading to an tight upper bound on the tidal torque. Since this cannot be explained by previous theoretical predictions for non-synchronized star-planet systems, our result requires mechanisms that suppress the tidal interaction in such systems.
@@ 1 Introduction
One of the representative examples that are not common in the Solar system is close-in Jovian planets. The firstly discovered exoplanet was such a “hot Jupiter”, 51 Pegasi b orbiting a solar-like star (Mayor & Queloz, 1995), and since then, hot Jupiters have been observed at a high frequency, whereas there is a caveat that hot Jupiters are preferentially identified by the radial velocity and planetary transit techniques. They are commonly considered to have formed beyond the ice line and migrated inward to their current positions (see, e.g., Dawson & Johnson, 2018; Fortney et al., 2021). The migration of planets is induced by interactions with the central star, disk, and other planets/companions (see, e.g., Perryman, 2018, Section 10.10 and 10.11).
In particular, the tidal interaction between a low-mass star and a close-in Jovian exoplanet is one of the key mechanisms of the star-planet interaction. The tidal dissipation in the stellar interior, which we focus on hereafter, plays an important role in the last phase of planetary migration. Specifically, the tidal interaction transports the orbital angular momentum (AM) to the stellar spin, reducing the semi-major axis, and obliquity of the planetary orbit (Darwin, 1880; Zahn, 1977; Hut, 1981; Barker & Ogilvie, 2009). While the orbital decay is confirmed for only one system, WASP-12 b (e.g., Yee et al., 2020), low-obliquity trends in close-in companions orbiting a low-mass star are acquired statistically (Winn et al., 2010; Albrecht et al., 2012, 2022). Also, the AM transport causes a spin-up of the host star, and this tendency is observationally obtained by the statistical examination (Tejada Arevalo et al., 2021). In addition to the spin-up effect, the tidal interaction has also attracted attention in terms of stellar magnetism because it may increase stellar magnetic activity (Cuntz et al., 2000). However, despite the importance, quantitatively consistent treatment of the efficiency of the tidal interaction has yet to be established because of both theoretical and observational difficulties.
@@
The main purpose of this paper is to investigate the method to restrict the efficiency of tidal interaction without facing the above uncertainties. Then, we focus on the non-synchronized system composed of a low-mass star and a close-in Jovian planet in order to simplify the physical picture. Also, considering the differential equations governing the AM co-evolution from a mathematical viewpoint of the dynamical system, we propose a novel method to impose a constraint on the efficiency of tidal interaction. A great advantage of our method is that we require only minimal assumptions about the model of tidal interaction and about the age and initial condition of the system, respectively.
@@ 2.1.3 Formulation of tidal torque
In a traditional framework (e.g., Goldreich & Soter, 1966), the efficiency of tidal interaction is characterized by the modified (dimensionless) quality factor $Q^{\prime}$, which is the inverse of the phase lag between the tidal potential and the tidal bulge divided by the tidal Love number (see Ogilvie, 2014). The tidal interaction is formulated using $Q^{\prime}$ as (see, e.g., Murray & Dermott, 1999; Benbakoura et al., 2019)
@@
Determining the value of $Q^{\prime}$ has remained a complex problem in astrophysics for a long time. This is because, as mentioned in Section 1, the tidal interaction originates from the combination of various non-linear mechanisms: the equilibrium/non-wave-like tide (Zahn, 1966; Zahn & Bouchet, 1989; Remus et al., 2012; Mathis et al., 2016), the dynamical tide in a convective zone induced by inertial wave (Ogilvie & Lin, 2007; Ogilvie, 2013; Mathis, 2015) and the dynamical tide in a radiative zone by internal gravity wave (Zahn, 1975; Goodman & Dickson, 1998; Barker & Ogilvie, 2010; Essick & Weinberg, 2016; Ma & Fuller, 2021). Given the controversial nature of determining the dominant mechanisms, we attempt to treat them comprehensively. For simplification, we assume the following two assumptions.
First, we exclude synchronized systems from our observational targets because it is not clear how tidal forces work on such systems. Then, we only consider the systems that satisfy $\Omega_{\mathrm{spin}}\ll\Omega_{\mathrm{orb}}$. In this case, since $\Gamma_{\mathrm{tide}}$ is positive, the system is subject to the orbital decay at the observed moment (see equations 6 and 9). We note that previous theoretical models often assume this condition for simplification (e.g., Barker, 2020), and hence, we can directly compare our results with theirs.
Second, we assume that $Q^{\prime}$ is approximated by the power of the tidal frequency, $\omega_{\mathrm{tide}}\equiv 2|\Omega_{\mathrm{orb}}-\Omega_{\mathrm{spin}}|$, with a floor value:
$Q^{\prime}=\max\left(Q^{\prime}_{\mathrm{0}}\left(\frac{\omega_{\mathrm{tide}}/2\pi}{2\,\mathrm{d^{-1}}}\right)^{-q},Q^{\prime}_{\mathrm{min}}\right),$ (10)
@@
The power-law relation in equation (10) comes from theoretical suggestion by previous papers dealing with the tide in the radiative zone of low-mass stars (Barker & Ogilvie, 2010; Essick & Weinberg, 2016; Barker, 2020; Ma & Fuller, 2021). We note that Barker (2020) claims that the tide in the radiative zone is dominant when $\Omega_{\mathrm{spin}}\ll\Omega_{\mathrm{orb}}$. However, as the detailed mechanism of tides is not still well understood (Appendix C), there are large uncertainties in $Q^{\prime}_{0}$ and $q$, which we treat as the free parameters to constrain $Q^{\prime}$.
The setting of the lower bound of $Q^{\prime}_{\mathrm{min}}$ in equation (10) is required to avoid the divergence of $\Gamma_{\mathrm{tide}}$ for $Q^{\prime}=0$ (equation 9) at the co-rotation radius where $\omega_{\mathrm{tide}}=0$ for a negative value of $q$. A non-zero $Q^{\prime}_{\mathrm{min}}$ is also reinforced from an observational point of view because there have been already a number of non-synchronized or eccentric close-in Jovian planets identified (e.g., Matsumura et al., 2008); these observations contradict a spurious enhancement in $\Gamma_{\mathrm{tide}}$ at the co-rotation radius because it makes close-in planets synchronized and circularized.
The presence of a lower bound of $Q^{\prime}$ is also required owing to the fact that a variation in the transit timing has not been detected for close-in planets (e.g., Ivshina & Winn, 2022), except for WASP-12 b (Yee et al., 2020; Turner et al., 2021; Wong et al., 2022). We note that the host star WASP-12 has $\sim 1.4\,M_{\odot}$ (Collins et al., 2017), which is larger than the low-mass stars we focus on.
@@ 1 $+$ -- -- --
Figure 3: Schematic pictures showing the reason for the existence of the forbidden area. The red shaded region and orange dotted arrows correspond to the initial condition and trajectories, respectively, which are determined by astrophysical conditions. The dark blue line represents the unstable manifold, which is one of the solutions of equation (11) passing through the origin. Because a mathematical argument shows that the unstable manifold can not be intersected (the black arrow with comment), we can show the existence of the forbidden region (the light blue hatched region).
@@
We calculate the evolution of $(\Omega_{\mathrm{orb}},\Omega_{\mathrm{spin}})$ for standard values of the wind torque parameter, $\alpha_{\mathrm{mb}}=1.5\times 10^{31}\,\mathrm{erg}$, reproducing the observed spin-down trend (see Appendix B) and the tidal parameters, $q=8/3$ and $Q^{\prime}_{0}=10^{5.5}$, adopted from the wave-breaking model (Barker & Ogilvie, 2010; Barker, 2020, see Appendix C), whereas the qualitative behavior described below is preserved regardless of the choice of these parameters. We also assume that the stellar properties, including the moment of inertia, $I_{\mathrm{s}}=0.0434\,M_{\odot}R_{\odot}^{2}$, determined from the stellar evolution code mesa (r12778; Paxton et al., 2011, 2013, 2015, 2018, 2019, see also Appendix D), are unchanged with time.
Figure 2 shows trajectories in the phase space (the red lines) for the initial stellar spin, $\Omega_{\mathrm{spin,init}}/2\pi=1.0\,\mathrm{d^{-1}}$. The trajectories can be separated into three evolutionary stages. In the first stage, the stellar spin is decelerated both through the wind torque and the tidal torque because $\Omega_{\mathrm{spin}}>\Omega_{\mathrm{orb}}$ (see also Figure 1(a)). In the next stage, the tidal torque begins to transport AM from the planet to the star as $\Omega_{\mathrm{spin}}<\Omega_{\mathrm{orb}}$ (see also Figure 1(b)), while a star continues to spin down because $\Gamma_{\mathrm{wind}}>\Gamma_{\mathrm{tide}}$. In the final stage, the tidal torque overcomes the wind torque, and then both stellar spin and planetary revolution are accelerated (stage 3). This tendency is summarized in Table 2.
A characteristic feature of these trajectories is that they never cross a particular curve (the black dashed line). That is to say, there is the forbidden area (the hatched area) in the phase space of ($\Omega_{\mathrm{orb}},\Omega_{\mathrm{spin}}$). The existence and property of the forbidden area are explained by combining astrophysical conditions and mathematical properties of the dynamical system theory. We proceed with the discussion without mathematical rigorous proofs used here, while a brief introduction is presented in Appendix E (see, e.g., Arnold, 2006; Hirsch et al., 2012, for more detail).
@@ 2.3 Procedure to constrain tidal quality factor
The novelty of our method is that the unstable manifold does not depend on time and the initial condition of the system. If one can fix the five “given parameters” listed in Table 1, a single set of observed $(\Omega_{\mathrm{orb}},\Omega_{\mathrm{spin}})$ in a star-planet system at any evolutionary stage can constrain the tidal quality factor in the two-dimensional $(q,Q^{\prime}_{0})$ parameter space. From an astrophysical point of view, our method gives an upper bound on the tidal torque from a close-in Jovian planet orbiting around a slowly rotating low-mass star. If such s system is observed in the lower right side of the phase space, it will provide a stringent constraint on the tidal torque because sufficiently weak tidal interaction is required to avoid the stellar spin-up by the transfer of AM from the planet. Since this methodology is simple and straightforward, our method is applicable to general tidal torque recipes (e.g., Barker, 2020) besides the parametrization in equation (10).
@@ 3 Target Selection
$I_{\mathrm{s}}$ $0.0434\,M_{\odot}R_{\odot}^{2}$2
@@
In this section, we briefly introduce the observational data we utilize. Following the argument in the previous section, our targets are non-synchronized systems consisting of a slowly rotating low-mass star and a close-in gaseous planet with a circular orbit (Section 2). Then, from the list of observed planetary systems of NASA Exoplanet Archive [^6] https://exoplanetarchive.ipac.caltech.edu/ (Akeson et al., 2013), we extract the stars satisfying the following conditions: $M_{\mathrm{s}}\leq 0.8\,M_{\odot}$, $M_{\mathrm{p}}\geq 0.5\,M_{\mathrm{J}}$, $0.02\leq\Omega_{\mathrm{spin}}/2\pi\leq 0.1\,\mathrm{d^{-1}}$, $\Omega_{\mathrm{orb}}/2\pi\geq 0.2\,\mathrm{d^{-1}}$, and an eccentricity $\leq 0.1$. Figure 5 shows the plot of the extracted samples (colored symbols) in the phase space of ($\Omega_{\mathrm{orb}},\Omega_{\mathrm{spin}}$) among other samples satisfying $M_{\mathrm{p}}\geq 0.5\,M_{\mathrm{J}}$ (small gray symbols). We can confirm that no data is distributed in the lower-right region of $\Omega_{\mathrm{orb}}$$\Omega_{\mathrm{spin}}$ plane; the shape of the empty region looks exactly like a forbidden region in Figures 3 and 4.

### v2

@@ Abstract
We propose a novel method to constrain the tidal quality factor, $Q^{\prime}$, from an observed non-synchronized star-planet system consisting of a slowly rotating low-mass star and a close-in Jovian planet, taking into account the co-evolution of stellar spin and planetary orbit due to the tidal interaction and the magnetic braking. On the basis of dynamical system theory, the track of the co-evolution of angular momentum from the fast rotator regime for such a system exhibits the existence of a forbidden region in the $\Omega_{\mathrm{orb}}$$\Omega_{\mathrm{spin}}$ plane , where $\Omega_{\mathrm{spin}}$ and $\Omega_{\mathrm{orb}}$ denote the angular velocity of the stellar spin and planetary orbit, respectively. The forbidden region is determined primarily by the strength of the tidal interaction. By comparing ($\Omega_{\mathrm{orb}},\Omega_{\mathrm{spin}}$) of a single star-planet system to the forbidden region, we can constrain the tidal quality factor regardless of the evolutionary history of the system. The application of this method to the star-planet system, NGTS-10 – NGTS-10 b, gives $Q^{\prime}\gtrsim 10^{8}$, leading to an tight upper bound on the tidal torque. Since this cannot be explained by previous theoretical predictions for non-synchronized star-planet systems, our result requires mechanisms that suppress the tidal interaction in such systems.
@@ 1 Introduction
One of the representative examples that are not common in the Solar system is close-in Jovian planets. The firstly discovered exoplanet orbiting a solar-like star was such a “hot Jupiter”, 51 Pegasi b (Mayor & Queloz, 1995), and since then, hot Jupiters have been observed at a high frequency, whereas there is a caveat that hot Jupiters are preferentially identified by the radial velocity and planetary transit techniques. They are commonly considered to have formed beyond the ice line and migrated inward to their current positions (see, e.g., Dawson & Johnson, 2018; Fortney et al., 2021). The migration of planets is induced by interactions with the central star, disk, and other planets/companions (see, e.g., Perryman, 2018, Section 10.10 and 10.11).
In particular, the tidal interaction between a low-mass star and a close-in Jovian exoplanet is one of the key mechanisms of the star-planet interaction. The tidal dissipation in the stellar interior, which we focus on hereafter, plays an important role in the last phase of planetary migration. Specifically, when the stellar spin is slow, the tidal interaction transports the orbital angular momentum (AM) to the stellar spin, reducing the semi-major axis and spin-orbit inclination angle of the planetary orbit (Darwin, 1880; Zahn, 1977; Hut, 1981; Barker & Ogilvie, 2009). While the orbital decay is confirmed for only one system, WASP-12 b (e.g., Yee et al., 2020), low-inclination trends in close-in companions orbiting around a low-mass star are acquired statistically (Winn et al., 2010; Albrecht et al., 2012, 2022). Also, the AM transport causes a spin-up of the host star, and this tendency is observationally obtained by the statistical examination (Tejada Arevalo et al., 2021). In addition to the spin-up effect, the tidal interaction has also attracted attention in terms of stellar magnetism because it may increase stellar magnetic activity (Cuntz et al., 2000). However, despite the importance, quantitatively consistent treatment of the efficiency of the tidal interaction has yet to be established because of both theoretical and observational difficulties.
@@
The main purpose of this paper is to investigate the method to restrict the efficiency of tidal interaction without facing the above uncertainties. Then, we focus on the non-synchronized system composed of a slowly-rotating low-mass star and a close-in Jovian planet in order to simplify the physical picture. Also, considering the differential equations governing the AM co-evolution from a mathematical viewpoint of the dynamical system, we propose a novel method to impose a constraint on the efficiency of tidal interaction. A great advantage of our method is that we require only minimal assumptions about the model of tidal interaction and about the age and initial condition of the system, respectively.
@@ 2.1.3 Formulation of tidal torque
In a traditional framework (e.g., Goldreich & Soter, 1966), the efficiency of tidal interaction is characterized by the modified (dimensionless) quality factor $Q^{\prime}$, which is inversely proportional to the phase lag between the tidal potential and the tidal bulge divided by the tidal Love number (see Ogilvie, 2014). The tidal interaction is formulated using $Q^{\prime}$ as (see, e.g., Murray & Dermott, 1999; Benbakoura et al., 2019)
@@
Determining the specific form of $Q^{\prime}$ remains a complex problem. Here, we assume that $Q^{\prime}$ is approximated by the power of the tidal frequency, $\omega_{\mathrm{tide}}\equiv 2|\Omega_{\mathrm{orb}}-\Omega_{\mathrm{spin}}|$, with a floor value:
$Q^{\prime}=\max\left[Q^{\prime}_{\mathrm{0}}\left(\frac{\omega_{\mathrm{tide}}/2\pi}{2\,\mathrm{d^{-1}}}\right)^{-q},Q^{\prime}_{\mathrm{min}}\right],$ (10)
@@
The power-law relation in equation (10) comes from theoretical suggestion by previous papers dealing with the tide in the radiative zone of low-mass stars induced by internal gravity wave (Zahn, 1975; Goodman & Dickson, 1998; Barker & Ogilvie, 2010; Essick & Weinberg, 2016; Ma & Fuller, 2021). We note that, while the equilibrium/non-wave-like tide (Zahn, 1966; Zahn & Bouchet, 1989; Remus et al., 2012; Mathis et al., 2016) and the dynamical tide in a convective zone induced by inertial wave (Ogilvie & Lin, 2007; Ogilvie, 2013; Mathis, 2015) are generally suggested as notable mechanisms, Barker (2020) claims that they are subdominant for slowly rotating solar-type stars, which are the main target of this paper (see Sections 2.3). However, even when limited to the tide in the radiative zone, the detailed mechanism is not still well understood (Appendix C) and there are large uncertainties in $Q^{\prime}_{0}$ and $q$. Then, we treat them as the free parameters to constrain $Q^{\prime}$.
The setting of the lower bound of $Q^{\prime}_{\mathrm{min}}$ in equation (10) is required to avoid the divergence of $\Gamma_{\mathrm{tide}}(\propto 1/Q^{\prime})$ for $Q^{\prime}=0$ (equation 9) at the co-rotation radius where $\omega_{\mathrm{tide}}=0$ for a negative value of $q$ (cf. Penev et al., 2018). A non-zero $Q^{\prime}_{\mathrm{min}}$ is also reinforced from an observational point of view because there have been already a number of non-synchronized or eccentric close-in Jovian planets identified (e.g., Matsumura et al., 2008, 2010); these observations contradict a spurious enhancement in $\Gamma_{\mathrm{tide}}$ at the co-rotation radius because it makes close-in planets synchronized and circularized. We note that the divergence of $Q^{\prime}$ at $\omega_{\mathrm{tide}}=0$ does not cause problematic behavior for positive $q$ because it simply gives $\Gamma_{\mathrm{tide}}(\propto 1/Q^{\prime})=0$.
The presence of a lower bound of $Q^{\prime}$ is also required owing to the fact that a variation in the transit timing has not been detected for close-in planets (e.g., Ivshina & Winn, 2022; Adams et al., 2024), except for WASP-12 b (Yee et al., 2020; Turner et al., 2021; Wong et al., 2022). We note that the host star WASP-12 has $\sim 1.4\,M_{\odot}$ (Collins et al., 2017), which is larger than the low-mass stars we focus on.
@@ 1 $+$ -- -- --
Figure 3: Schematic pictures showing the reason for the existence of the forbidden area. The dark green shaded region and light green dotted arrows correspond to the initial condition and trajectories, respectively, which are determined by astrophysical conditions. The dark blue line represents the unstable manifold, which is one of the solutions of equation (11) passing through the origin. Because a mathematical argument shows that the unstable manifold can not be intersected (the black arrow with comment), we can show the existence of the forbidden region (the light blue hatched region).
@@
We calculate the evolution of $(\Omega_{\mathrm{orb}},\Omega_{\mathrm{spin}})$ for standard values of the wind torque parameter, $\alpha_{\mathrm{mb}}=1.5\times 10^{31}\,\mathrm{erg}$, reproducing the observed spin-down trend (see Appendix B) and the tidal parameters, $q=8/3$ and $Q^{\prime}_{0}=10^{5.5}$, adopted from the wave-breaking model (Barker & Ogilvie, 2010; Barker, 2020, see Appendix C), whereas the qualitative behavior described below is preserved regardless of the choice of these parameters. We also assume that the stellar properties, including the moment of inertia, $I_{\mathrm{s}}=0.0435\,M_{\odot}R_{\odot}^{2}$, determined from the stellar evolution code mesa (r12778; Paxton et al., 2011, 2013, 2015, 2018, 2019, see also Appendix D), are unchanged with time.
Figure 2 shows trajectories in the phase space (the red lines) for the initial stellar spin, $\Omega_{\mathrm{spin,init}}/2\pi=1.0\,\mathrm{d^{-1}}$. The trajectories can be separated into three evolutionary stages, listed in Table 2. First, in the stage 1, the stellar spin is decelerated both through the wind torque and the tidal torque because $\Omega_{\mathrm{spin}}>\Omega_{\mathrm{orb}}$ (see also Figure 1(a)). While $\Omega_{\mathrm{orb}}$ also decreases in the stage 1, it is not explicitly visible in Figure 2 because $\Gamma_{\mathrm{wind}}\gg\Gamma_{\mathrm{tide}}$. Next, in the stage 2, the tidal torque begins to transport AM from the planet to the star as $\Omega_{\mathrm{spin}}<\Omega_{\mathrm{orb}}$ (see also Figure 1(b)), while a star continues to spin down because $\Gamma_{\mathrm{wind}}>\Gamma_{\mathrm{tide}}$. Finally, in the stage 3, the tidal torque overcomes the wind torque, and then both stellar spin and planetary revolution are accelerated. This tendency is summarized in Table 3. We note that the observational targets we utilize to constrain the tidal torque are the systems including slowly rotating stars in the stages 2 and 3 listed in Table 2.
A characteristic feature of these trajectories is that they never cross a particular curve (the black dashed line). That is to say, there is the forbidden area (the hatched area) in the phase space of ($\Omega_{\mathrm{orb}},\Omega_{\mathrm{spin}}$). The existence and property of the forbidden area are explained by combining astrophysical conditions and mathematical properties of the dynamical system theory. We proceed with the discussion without mathematical rigorous proofs used here, while a brief introduction is presented in Appendix E (see, e.g., Arnold, 1992; Hsieh & Sibuya, 1999, for more detail).
@@ 2.3 Procedure to constrain tidal quality factor
The novelty of our method is that the unstable manifold does not depend on time and the initial condition of the system. If one can fix the five “given parameters” listed in Table 1, a single set of observed $(\Omega_{\mathrm{orb}},\Omega_{\mathrm{spin}})$ in a star-planet system at any evolutionary stage can constrain the tidal quality factor in the two-dimensional $(q,Q^{\prime}_{0})$ parameter space. Although this study does not take into account finite eccentricity, $e$, and spin-orbit inclination, $i$, our method can be applied without error as long as $e$ and $i$ are (nearly) zero at the time of observation, even if these took finite values during the evolution before the observation (see the discussion in Section 5.3).
From an astrophysical point of view, our method gives an upper bound on the tidal torque from a close-in Jovian planet orbiting around a slowly rotating low-mass star. If such s system is observed in the lower right side of the phase space, it will provide a stringent constraint on the tidal torque because sufficiently weak tidal interaction is required to avoid the stellar spin-up by the transfer of AM from the planet. Since this methodology is simple and straightforward, our method is applicable to general tidal torque recipes (e.g., Barker, 2020) besides the parametrization in equation (10).
@@ 3 Target Selection
$I_{\mathrm{s}}$ $0.0435\,M_{\odot}R_{\odot}^{2}$2
@@
In this section, we briefly introduce the observational data we utilize. Our targets are non-synchronized systems consisting of a slowly rotating low-mass star and a close-in gaseous planet with a circular orbit because their observation can be directly compared with our model (Section 2). Then, from the list of observed planetary systems of NASA Exoplanet Archive [^6] https://exoplanetarchive.ipac.caltech.edu/ (Akeson et al., 2013), we extract the stars satisfying the following conditions: $M_{\mathrm{s}}\leq 0.8\,M_{\odot}$, $M_{\mathrm{p}}\geq 0.5\,M_{\mathrm{J}}$, $0.02\leq\Omega_{\mathrm{spin}}/2\pi\leq 0.1\,\mathrm{d^{-1}}$, $\Omega_{\mathrm{orb}}/2\pi\geq 0.2\,\mathrm{d^{-1}}$, and $e\leq 0.1$. Figure 5 shows the plot of the extracted samples (colored symbols) in the phase space of ($\Omega_{\mathrm{orb}},\Omega_{\mathrm{spin}}$) among other samples satisfying $M_{\mathrm{p}}\geq 0.5\,M_{\mathrm{J}}$ (small gray symbols). We can confirm that no data is distributed in the lower-right region of $\Omega_{\mathrm{orb}}$$\Omega_{\mathrm{spin}}$ plane; the shape of the empty region looks exactly like a forbidden region in Figures 3 and 4.